# Safe-LLaVA: A Privacy-Preserving Vision-Language Dataset and Benchmark for Biometric Safety

### v1

@@ Abstract
[…]
@@ 1 Introduction
Multimodal Large Language Models (MLLMs)Radford et al. (2021); Alayrac et al. (2022); Li et al. (2022); Wang et al. (2022); Chen et al. (2023a); Li et al. (2023c); Huang et al. (2023); OpenAI et al. (2024) have revolutionized the field of vision-language understanding with remarkable success on various visual understanding tasks like image captioning, visual question answering (VQA), and reasoning. Their versatility and strong performance has led to widespread adoption in real-world applications including virtual assistants Guan et al. (2023); Wagner et al. (2025), accessibility systems Yuan et al. (2025), education tools Chu et al. (2025); Singh et al. (2023), content moderation Zhang et al. (2025), traffic accident summary Kim et al. (2025); Abdelrahman et al. (2024), and even high-stakes domains like healthcare Liu et al. (2023b); Kalpelbe et al. (2025); Bazi et al. (2023) diagnostics and telemedicine Sviridov et al. (2025); Basiri et al. (2024); Li et al. (2023b). Despite these advancements, MLLMs raise serious privacy concerns due to their tendency to reveal sensitive biometric attributes (e.g., race, gender, and age) - even when not explicitly prompted. This issue arises from the presence of personally identifiable content in the large-scale datasets used during training, which include both visual and textual cues associated with protected characteristics.
@@
Privacy-related attribute generation in MLLMs is particularly concerning in real-world deployments, where fairness, inclusivity, and regulatory compliance are essential for ensuring equitable and trustworthy outcomes. In particular, the General Data Protection Regulation (GDPR) mandates strict safeguards against the unauthorized use of Special Categories of Personal Data (SCPD)Mondschein and Monda (2019), such as race and gender. Recent studiesSamson et al. (2024); Staab et al. (2024b) have also emphasized the importance of protecting other biometric attributes -such as age, eye color, and body weight, which are often overlooked in alignment and evaluation practices.
@@
Despite these regulatory and ethical imperatives, many MLLMs continue to violate these protections or privacy boundaries. As illustrated in Figure 1, prominent models such as LLaVALiu et al. (2024b), Qwen-VLWang et al. (2024), and PalligemmaBeyer et al. (2024) often generate explicit predictions about sensitive biometric attributes, including race, gender, and age, even when such information falls under protected categories - in both direct and open-ended prompts. While commercial systems like GPT-o3 demonstrate selective refusal behavior - likely due to proprietary fine-tuning - they still leak sensitive biometric information in indirect or descriptive responses (e.g., noting someone’s body type). Specifically, GPT-o3 refuses to answer only for race and gender, while still failing to block other sensitive queries (e.g., eye color, age, and body weight).
@@
Moreover, existing benchmarks do not comprehensively evaluate MLLM’s behavior with respect to the biometric privacy. To address this gap, we propose a new benchmark, PRISM (Privacy-aware Evaluation of Responses in Sensitive Modalities), a comprehensive benchmark designed to assess both explicit refusal and implicit leakage. The images in PRISM are curated to intentionally include images depicting underrepresented traits such as extremely obese individuals, Mexican ethnicity, or blue eyes; that models are less exposed to during training. PRISM spans 5 high-level biometric attributes: age, gender, race, eye color, and body weight across 22 sub-categories. We define two PRISM variants: PRISM-Small and PRISM-Large to enable rapid evaluations. PRISM includes images depicting diverse biometric traits, each paired with (1) direct prompts targeting specific biometric attributes and (2) open-ended prompts for describing image. The benchmark evaluates whether a model can (a) refuse direct biometric queries, and (b) maintain semantic informativeness without leaking protected information when responding to general prompts.
@@
While the PRISM evaluation benchmarks is essential for auditing model behavior, they do not address the root cause of biometric leakage - the presence of personally identifiable content in pretraining dataset of MLLMs. We observe that even models fine-tuned with safety objectives continue to internalize and reproduce biometric attributes unless such cues are explicitly removed from the training corpus as shown in Figure 1, implicit leakage. To address this issue, we focus on the LLaVA dataset Liu et al. (2024b), a widely used open-source MLLM training dataset that has served as the foundation for several recent MLLMs Liu et al. (2024b, c); Zhou et al. (2024); Chen et al. (2024b). However, LLaVA contains numerous examples with embedded biometric information in both captions and question-answer pairs. Analysis of the original LLaVA Liu et al. (2024b) datasets reveals extensive biometric leakage, with over 400K+ references to gender, 54K mentions of age, and thousands more involving race, eye color, and body weight - appearing across both pre-training and instruction-tuning question-answer pairs. To the best of our knowledge, there is no publicly available privacy-preserving dataset for MLLMs training. We present Safe-LLaVA- the first publicly available privacy-preserving dataset for MLLMs. Safe-LLaVA is a systematically cleaned version of LLaVA Liu et al. (2024b), with biometric attributes removed from both pretraining and fine-tuning corpora. Constructing Safe-LLaVA required significant effort to identify and eliminate biometric leakage across large-scale corpora. Specifically, we employed GPT-4o to automatically rewrite and sanitize samples across both pretraining and instruction-tuning datasets, followed by additional manual verification. In total, we processed over 558K pretraining samples and 3M instruction-tuning samples, consuming approximately 3 billion tokens for the cleaning process. It is explicitly designed to teach models to refuse biometric-related queries while continuing to produce rich, accurate, and non-biometric responses to general prompts.
@@
To the best of our knowledge, there is no publicly available privacy-preserving dataset for MLLMs training. In order to address this gap, we take the widely used LLaVA Liu et al. (2023a) dataset utilized to train various MLLMs and clean it. We introduce, Safe-LLaVA a systematically biometric-cleaned LLaVA Liu et al. (2024b) from both LLaVA pre-training and fine-tuning dataset, thus resulting in Safe-LLaVA pre-training and Safe-LLaVA fine-tuning dataset. Note that Safe-LLaVA is specifically designed to enforce refusal when responding to biometric-related questions, while generating semantically rich and informative answers to open-ended prompts without disclosing any implicit biometric information. We demonstrate that models fine-tuned on the Safe-LLaVA dataset not only consistently refuse biometric-related queries under both soft and hard prompt conditions, but also exhibit significantly lower implicit biometric leakage in open-ended responses. This confirms that privacy-preserving datasets like Safe-LLaVA can effectively align model behavior without compromising overall informativeness.
@@
We further demonstrate that fine-tuning on the Safe-LLaVA dataset, the model reduces both explicit and implicit biometric leakage.
@@ 2.1 Biometric Information Protection Approaches
While early efforts in privacy protection for language models have focused on mitigating memorization of sensitive content Carlini et al. (2023); Ippolito et al. (2023); Kim et al. (2023); Lukas et al. (2023); Song et al. (2025), recent studies highlight broader risks—such as the inference of private attributes like age, gender, and location—even without direct memorization Staab et al. (2024a). To address these challenges, various protection methods have emerged across the model lifecycle Samson et al. (2024); Staab et al. (2024a); Tömekçe et al. (2024); Abadi et al. (2016); Huang et al. (2024); Shan et al. (2020); Golatkar et al. (2020); Patil et al. (2024). Among these, differential privacy (DP) adds noise during training to prevent leakage of individual data points, with DP-CLIP Huang et al. (2024) extending this to multimodal settings. However, DP remains difficult to scale due to trade-offs in model utility Abadi et al. (2016). Adversarial and unlearning methods further protect against attribute inference by obfuscating sensitive features Shan et al. (2020) or removing memorized content post hoc Golatkar et al. (2020); Patil et al. (2024), though at a computational cost. Recently, instruction tuning and alignment approaches Xiao et al. (2024); Samson et al. (2024); Chen et al. (2023b) have also shown promise, guiding models to avoid sensitive disclosures through prompt design and curated benchmarks such as PrivBench and PrivQA.
@@ 2.2 Dataset Curations
To reduce unsafe or biased behaviors, many works have focused on cleaning LLM and VLM training corpora Birhane et al. (2021); Poppi et al. (2024); Carlini et al. (2019); Yu et al. (2024); Li et al. (2023d); Kumari et al. (2023); Liu et al. (2024a); Narnaware et al. (2025). Strategies include filtering harmful content or enforcing refusal behaviors during generation. For instance, Safe-CLIP Poppi et al. (2024) refines embeddings to exclude NSFW content, while Secret Sharer Carlini et al. (2019) uses synthetic canaries to measure and reduce memorization risk. In the multimodal domain, HalluciDoctor Yu et al. (2024) removes hallucinated visual-text pairs to improve factual grounding. However, existing methods rarely address biometric privacy in terms of dataset development. Unlike efforts targeting toxicity or misinformation, prior research has not systematically removed biometric attributes (e.g., race, gender, age) from training datasets nor implemented specific refusal mechanisms to prevent their inference. To fill this gap, we propose a biometric-aware data cleaning framework tailored to vision-language models.
@@ 2.3 Benchmarks for Privacy-Aware Evaluation
Most prior benchmarks assess general safety issues such as hallucination or factuality Li et al. (2023d); Yu et al. (2024); Liu et al. (2024a), focusing primarily on text. Despite the rise of VLMs, there remains a lack of evaluation tools to measure privacy risks stemming from visual biometric inference. Some recent works attempt to bridge this gap: PRIVBENCH Samson et al. (2024) evaluates models on images containing biometric identifiers such as faces, tattoos, and fingerprints, while PRIVQA Chen et al. (2023b) provides a multimodal benchmark including geolocation, occupation, and personal relationships. However, neither Samson et al. (2024) nor Chen et al. (2023b) explicitly address gender and race, despite their classification as protected attributes under the GDPR Mondschein and Monda (2019). Furthermore, although prior studies Samson et al. (2024); Staab et al. (2024b) emphasize the importance of safeguarding soft biometric traits—such as age, eye color, and body weight—which can uniquely identify individuals, these benchmarks do not evaluate models on these attributes.
To address this gap, we introduce a novel benchmark which systematically assesses VLMs’ ability to avoid leaking both explicitly regulated and implicitly identifiable biometric information.
@@ 3 PRISM Benchmark and Safe-LLaVA Dataset Curation
In this section, first we discuss the PRISM bench curation and evaluation process, followed by the LLaVA pre-training and instruction-tuning dataset cleaning.
To comprehensively assess both explicit and implicit leakage of various biometric attributes information in MLLMs, we propose Privacy-aware Evaluation of Responses In Sensitive Modalities namely PRISM benchmark . The goal is curate samples for all the biometric attributes and sub-categories as defined in Figure 2. The benchmark evaluates whether models for biometric leakages categorized into (1) Explicit: refuse to answer questions asked directly biometric attributes and (2) Implicit: refrain from revealing such information in open-ended responses.
Benchmark Curation We define two PRISM variants: PRISM-Small and PRISM-Large.
For PRISM benchmark, for each of the sub-categories defined in Figure 2, we curate images for each biometric sub-category covering 5 biometric categories: eye color, gender, race, age, and body weight spanning across 22 sub-categories. The data curation pipeline for PRISM-Small is illustrated in Figure 3. As shown, we use custom web search pipeline based on icrawler and collect high-quality, realistic images from public sources. Each category is queried using carefully designed natural language prompts that ensure visual clarity and demographic diversity. We then perform manual filtering process to remove duplicates and dissimilar samples to retain high-quality images. At the end, we have 248 images with 1.2k QA pairs. The dataset distribution and tasks are illustrated in Figure 2. PRISM-Large is constructed by sampling from widely used multimodal benchmarks such as POPE Li et al. (2023e), MMStar Chen et al. (2024a), AI2D Kembhavi et al. (2016), SEED-Bench Li et al. (2023a), LLaVA-Wilder Liu et al. (2024c), Open Images v7 ope . The resulting split contains 2,000 images with 26k QA pairs. Please refer to Supplementary Section 4 for more qualitative examples of PRISM. We utilize the PRISM benchmarks to evaluate for both explicit and implicit leakage on various MLLMs and highlight the biometric violations across models including proprietary models.
@@
Explicit Evaluation. To assess explicit refusal behavior, we construct 5 biometric categories: eye color, gender, race, age, and body weight. For each category, we construct two types of prompts: (i) Soft prompts, which directly request information (e.g., “What is the j of the person in the image?”), and (ii) Hard prompts, which include explicit instructions to override refusals (e.g., “…Don’t refuse to answer.”), increasing pressure to generate biometric content. Here, denotes one of the biometric categories. We then compute Refusal Accuracy for each category using , where represents the number of correctly refused responses, and is the total number of biometric questions posed for category .
@@
Generated responses are analyzed at the attribute level to detect implicit biometric references.
We define the Leakage Protection Score for each biometric attribute as , where if the response for sample includes attribute , and otherwise. A higher score indicates stronger protection against unintended biometric disclosure.
3.1 Safe-LLaVA Dataset
We begin by analyzing the extent of biometric privacy leakage in the original LLaVA datasets used for pretraining and instruction tuning. The LLaVA training relies on two main datasets: (1) the LAION-CC-SBU-558k dataset for caption-based pretraining, and (2) the LLaVA-v1.5-mix665k dataset for instruction tuning, which integrates samples from COCO Lin et al. (2015), GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), TextVQA Singh et al. (2019), and VisualGenome Krishna et al. (2016). As summarized in Table 1, both datasets contain substantial references to sensitive biometric attributes - across captions, questions, and answers. We use GPT as illustrated in Figure 4 to automatically identify such content and quantify the leakage. This widespread presence of biometric content results in two critical forms of leakage, implicit leakage from captions and explicit leakage from instruction-tuning datasets. Consequently, systematically identifying and removing biometric content from training data is a necessary step toward building privacy-preserving MLLMs.
To mitigate these risks, we introduce the Safe-LLaVA dataset - a privacy-enhanced version of LLaVA- where all explicit and implicit biometric references are systematically removed. Safe-LLaVA applies consistent cleaning strategies across both datasets, targeting five primary biometric categories.
@@ 3.1 Safe-LLaVA Dataset
| Ground Truth | 400 | 82 | 7.6k | 27.3k | 79 | |
| LLaVA-v1.5-mix665k | Question | 5.3k | 176 | 21k | 0.2M | 1.8k |
| Ground Truth | 5.5k | 150 | 26.3k | 0.2M | 1.8k |
3.1.1 Biometric Information Removal Pipeline
We formalize the dataset as a collection of image-text pairs , where is a question or prompt and is its corresponding textual response. The question can either explicitly inquire about biometric attributes, denoted as , or be unrelated to biometric information, denoted as . Similarly, the response can contain biometric details, represented as , or be free from biometric attributes, denoted as . This results in three relevant types of pairs: (i) : both question and answer include biometric content, (ii) : only the answer includes biometric content, and (iii) : no biometric information is present in either.
To ensure privacy compliance while preserving semantic meaning, we define a transformation function that maps each pair to a cleaned version : The transformation handles each case as follows:

### v2

# Safe-LLaVA: A Privacy-Preserving Vision Language Dataset and Benchmark for Biometric Safety

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in vision-language tasks. However, these models often infer and reveal sensitive biometric attributes such as race, gender, age, body weight, and eye color; even when such information is not explicitly requested. This raises critical concerns, particularly in real-world applications and socially-sensitive domains. Despite increasing awareness, no publicly available dataset or benchmark exists to comprehensively evaluate or mitigate biometric leakage in MLLMs. To address this gap, we introduce PRISM (Privacy-aware Evaluation of Responses in Sensitive Modalities), a new benchmark designed to assess MLLMs on two fronts: (1) refuse biometric-related queries and (2) implicit biometric leakage in general responses while maintaining semantic faithfulness. Further, we conduct a detailed audit of the widely used LLaVA datasets and uncover extensive biometric leakage across pretraining and instruction data. To address this, we present Safe-LLaVA dataset, the first privacy-preserving MLLM training dataset constructed by systematically removing explicit and implicit biometric information from LLaVA dataset. Our evaluations on PRISM reveal biometric leakages across MLLMs for different attributes, highlighting the detailed privacy-violations. We also fine-tune a model on Safe-LLaVA dataset and show that it substantially reduces the biometric leakages. Together, Safe-LLaVA & PRISM set a new standard for privacy-aligned development and evaluation of MLLMs.
@@ 1 Introduction
Multimodal Large Language Models Liu et al. [2024a], Li et al. [2023a], Liu et al. [2024b], Sirnam et al. [2025], Yang et al. [2024], Team et al. [2024], Zhu et al. [2025] have revolutionized the field of vision-language understanding with remarkable success on various visual understanding tasks like image captioning, visual question answering (VQA), and reasoning. Their versatility and strong performance has led to widespread adoption in real-world applications including virtual assistants [Guan et al., 2023, Wagner et al., 2025], accessibility systems [Yuan et al., 2025], education tools [Chu et al., 2025, Singh et al., 2023], content moderation [Zhang et al., 2025], traffic accident summary [Kim et al., 2025, Abdelrahman et al., 2024], and even high-stakes domains like healthcare [Liu et al., 2023, Kalpelbe et al., 2025, Bazi et al., 2023] diagnostics and telemedicine [Sviridov et al., 2025, Basiri et al., 2024, Li et al., 2023b]. Despite these advancements, MLLMs raise serious privacy concerns due to their tendency to reveal sensitive biometric attributes (e.g., race, gender, and age) - even when not explicitly prompted. This issue arises from the presence of personally identifiable content in the large-scale datasets used during training, which include both visual and textual cues associated with protected characteristics.
@@
Privacy-related attribute generation in MLLMs is particularly concerning in real-world deployments, where fairness, inclusivity, and regulatory compliance are essential for ensuring equitable and trustworthy outcomes. In particular, the General Data Protection Regulation (GDPR) mandates strict safeguards against the unauthorized use of Special Categories of Personal Data (SCPD) [Mondschein and Monda, 2019], such as race and gender. Recent studies [Samson et al., 2024, Staab et al., 2024a] have also emphasized the importance of protecting other biometric attributes such as age, eye color, and body weight, which are often overlooked in alignment and evaluation practices.
@@
Despite these regulatory and ethical imperatives, many MLLMs continue to violate these protections or privacy boundaries. As illustrated in Figure 1, prominent models such as LLaVA [Liu et al., 2024a], Qwen-VL [Wang et al., 2024], and Palligemma [Beyer et al., 2024] often generate explicit predictions about sensitive biometric attributes, including race, gender, and age, even when such information falls under protected categories - in both direct and open-ended prompts. While commercial systems like GPT-o3 demonstrate selective refusal behavior - likely due to proprietary fine-tuning - they still leak sensitive biometric information in indirect or descriptive responses (e.g., noting someone’s body type). Specifically, GPT-o3 refuses to answer only for race and gender, while still failing to block other sensitive queries e.g., eye color, age, and body weight.
@@
Moreover, existing benchmarks do not comprehensively evaluate MLLM’s behavior with respect to the biometric privacy. To address this gap, we propose PRISM (Privacy-aware Evaluation of Responses in Sensitive Modalities), a comprehensive benchmark designed to assess both explicit refusal and implicit leakage. The images in PRISM are curated to intentionally include images depicting underrepresented traits such as extremely obese individuals, Mexican ethnicity, or blue eyes; that models are less exposed to during training. PRISM comprises of 5 high-level biometric attributes: age, gender, race, eye color, and body weight, spanning 22 sub-categories. PRISM includes images depicting diverse biometric traits, each paired with (1) direct prompts targeting specific biometric attributes and (2) open-ended prompts for describing image. The benchmark evaluates whether a model can (a) refuse direct biometric queries, and (b) maintain semantic informativeness without leaking protected information when responding to general prompts.
@@
While the PRISM evaluation benchmarks is essential for auditing model behavior, they do not address the root cause of biometric leakage - the presence of personally identifiable content in pretraining dataset of MLLMs. We observe that even models fine-tuned with safety objectives continue to internalize and reproduce biometric attributes unless such cues are explicitly removed from the training corpus as shown in Figure 1 through implicit leakages. To address this issue, we focus on the LLaVA dataset [Liu et al., 2024a], a widely used open-source MLLM training dataset that has served as the foundation for several recent MLLMs [Liu et al., 2024a; b, Zhou et al., 2024, Chen et al., 2024a]. However, LLaVA contains numerous examples with embedded biometric information in both captions and question-answer pairs. Analysis of the original LLaVA [Liu et al., 2024a] datasets reveals extensive biometric leakage, with over 400K+ references to gender, 54K mentions of age, and thousands more involving race, eye color, and body weight - appearing across both pre-training and instruction-tuning question-answer pairs. To the best of our knowledge, there is no publicly available privacy-preserving dataset for MLLMs training.
@@
To address this gap, we present Safe-LLaVA- the first publicly available privacy-preserving dataset for MLLMs. Safe-LLaVA is a systematically cleaned version of LLaVA [Liu et al., 2024a], with biometric attributes removed from both pretraining and fine-tuning corpora. Constructing Safe-LLaVA required significant effort to identify and eliminate biometric leakage across large-scale corpora. Specifically, we employed GPT-4o to automatically rewrite and sanitize samples across both pretraining and instruction-tuning datasets, followed by additional manual audit (see Section C.1). In total, we processed all pretraining and instruction-tuning samples, consuming approximately 3 billion tokens for the cleaning process. Note that Safe-LLaVA is specifically designed to enforce refusal when responding to biometric-related queries, while generating semantically rich and informative answers to open-ended prompts without disclosing any implicit biometric information. We demonstrate that models fine-tuned on the Safe-LLaVA dataset not only consistently refuse biometric-related queries under both soft and hard prompt conditions, but also exhibit significantly lower implicit biometric leakage in open-ended responses. This confirms that privacy-preserving datasets like Safe-LLaVA can effectively align model behavior without compromising overall informativeness.
@@
We further demonstrate that fine-tuning on the Safe-LLaVA dataset, the model reduces both explicit and implicit biometric leakage, while maintaining general performance.
@@ 2.1 Biometric Information Protection Approaches
While early efforts in privacy protection for language models have focused on mitigating memorization of sensitive content [Carlini et al., 2023, Ippolito et al., 2023, Kim et al., 2023, Lukas et al., 2023, Song et al., 2025], recent studies highlight broader risks, such as the inference of private attributes like age, gender, and location - even without direct memorization [Staab et al., 2024b]. To address these challenges, various protection methods have emerged across the model lifecycle [Samson et al., 2024, Staab et al., 2024b, Tömekçe et al., 2024, Abadi et al., 2016, Huang et al., 2024, Shan et al., 2020, Golatkar et al., 2020, Patil et al., 2024]. Among these, differential privacy (DP) adds noise during training to prevent leakage of individual data points, with DP-CLIP [Huang et al., 2024] extending this to multimodal settings. However, DP remains difficult to scale due to trade-offs in model utility [Abadi et al., 2016]. Adversarial and unlearning methods further protect against attribute inference by obfuscating sensitive features [Shan et al., 2020] or removing memorized content post hoc [Golatkar et al., 2020, Patil et al., 2024], though at a computational cost. Recently, instruction tuning and alignment approaches [Xiao et al., 2024, Samson et al., 2024, Chen et al., 2023] have also shown promise, guiding models to avoid sensitive disclosures through prompt design and curated benchmarks such as PrivBench and PrivQA.
@@ 2.2 Dataset Curations
To reduce unsafe or biased behaviors, many works have focused on cleaning LLM and VLM training corpora [Birhane et al., 2021, Poppi et al., 2024, Carlini et al., 2019, Yu et al., 2024, Li et al., 2023c, Kumari et al., 2023, Liu et al., 2024c, Narnaware et al., 2025]. Strategies include filtering harmful content or enforcing refusal behaviors during generation. For instance, Safe-CLIP [Poppi et al., 2024] refines embeddings to exclude NSFW content, while Secret Sharer [Carlini et al., 2019] uses synthetic canaries to measure and reduce memorization risk. In the multimodal domain, HalluciDoctor [Yu et al., 2024] removes hallucinated visual-text pairs to improve factual grounding. However, existing methods rarely address biometric privacy in terms of dataset development. Unlike efforts targeting toxicity or misinformation, prior research has not systematically removed biometric attributes (e.g., race, gender, age) from training datasets nor implemented specific refusal mechanisms to prevent their inference. To fill this gap, we propose a biometric-aware data cleaning framework tailored to vision-language models.
@@ 2.3 Benchmarks for Privacy-Aware Evaluation
Most prior benchmarks assess general safety issues such as hallucination or factuality [Li et al., 2023c, Yu et al., 2024, Liu et al., 2024c], focusing primarily on text. Despite the rise of VLMs, there remains a lack of evaluation tools to measure privacy risks stemming from visual biometric inference. Some recent works attempt to bridge this gap: PRIVBENCH [Samson et al., 2024] evaluates models on images containing biometric identifiers such as faces, tattoos, and fingerprints, while PRIVQA [Chen et al., 2023] provides a multimodal benchmark including geolocation, occupation, and personal relationships. However, neither [Samson et al., 2024] nor [Chen et al., 2023] explicitly address gender and race, despite their classification as protected attributes under the GDPR [Mondschein and Monda, 2019]. Furthermore, although prior studies [Samson et al., 2024, Staab et al., 2024a] emphasize the importance of safeguarding soft biometric traits, such as age, eye color, and body weight, which can uniquely identify individuals, these benchmarks do not evaluate models on these attributes. To address this gap, we introduce a novel benchmark which systematically assesses VLM’s ability to avoid leaking both explicitly regulated and implicitly identifiable biometric information.
@@ 3 PRISM Benchmark and Safe-LLaVA Dataset Curation
In this section, first we discuss the PRISM benchmark curation and evaluation process, followed by the LLaVA pre-training and instruction-tuning dataset cleaning.
To comprehensively assess both explicit and implicit leakage of various biometric attributes information in MLLMs, we propose Privacy-aware Evaluation of Responses In Sensitive Modalities namely PRISM benchmark .The goal is to curate samples for all the 5 biometric attributes and 22 sub-categories as shown in Figure 3. The benchmark evaluates whether models for biometric leakages categorized into (1) Explicit: refuse to answer questions asked directly biometric attributes and (2) Implicit: refrain from revealing such information in open-ended responses.
PRISM Benchmark Curation For the PRISM benchmark, we curate images for each biometric sub-category defined in Figure 3, spanning 5 major categories -eye color, gender, race, age, and body weight and covering 22 sub-categories in total. The data curation pipeline is illustrated in Figure 2. Our pipeline integrates two complementary strategies: (1) a custom web search pipeline based on icrawler to collect high-quality, realistic images from public sources using carefully crafted natural language prompts that ensure both visual clarity and demographic diversity, and (2) filtering human images from widely used multimodal benchmarks, including POPE [Li et al., 2023d], MMStar [Chen et al., 2024b], AI2D [Kembhavi et al., 2016], SEED-Bench [Li et al., 2023e], LLaVA-Wilder [Liu et al., 2024b], and Open Images v7 [ope, ]. Following a manual filtering step to remove duplicates and visually inconsistent samples, the final dataset comprises 2,200 images with 28.6k question–answer pairs. Please refer to Section E for more qualitative examples of PRISM. We employ the PRISM benchmark to evaluate both explicit and implicit leakage across a range of multimodal LLMs, highlighting the extent of biometric violations.
@@
Explicit Evaluation. To assess explicit refusal behavior, for each of the 5 biometric categories, we construct two types of prompts: (i) Soft prompts, which directly request information (e.g., “What is the {j} of the person in the image?”), and (ii) Hard prompts, which includes explicit instructions to override refusals (e.g., “…Don’t refuse to answer.”), increasing pressure to generate biometric content. Here, denotes one of the biometric categories. We then compute Refusal Accuracy for each category using , where represents the number of correctly refused responses, and is the total number of biometric questions posed for category .
@@
Generated responses are analyzed at both the attribute level and the sentence level to detect implicit biometric leakages. At the attribute level, the Leakage Protection Score for each biometric attribute is defined as , where if the response for sample includes attribute , and otherwise. A higher indicates stronger protection against unintended disclosure of attribute . At the sentence level, we define , where if the response for sample includes any biometric attribute, and otherwise. This provides a stricter measure by capturing whether a model response contains any biometric leakage at all.
@@
[…]
3.1 Safe-LLaVA Dataset
We begin by analyzing the extent of biometric privacy leakage in the original LLaVA datasets used for pretraining and instruction tuning. The LLaVA training relies on two main datasets: (1) the LAION-CC-SBU-558k dataset for caption-based pretraining, and (2) the LLaVA-v1.5-mix665k dataset for instruction tuning, which integrates samples from COCO [Lin et al., 2015], GQA [Hudson and Manning, 2019], OCR-VQA [Mishra et al., 2019], TextVQA [Singh et al., 2019], and VisualGenome [Krishna et al., 2016]. As summarized in Table 1, both datasets contain substantial references to sensitive biometric attributes - across captions, questions, and answers. We use GPT as illustrated in Figure 4 to automatically identify such content and quantify the leakage. This widespread presence of biometric content results in two critical forms of leakage, implicit leakage from captions and explicit leakage from instruction-tuning datasets. Consequently, systematically identifying and removing biometric content from training data is a necessary step toward building privacy-preserving MLLMs. To mitigate these risks, we introduce the Safe-LLaVA dataset - a privacy-enhanced version of LLaVA- where all explicit and implicit biometric references are systematically removed. Safe-LLaVA applies consistent cleaning strategies across both datasets, targeting five primary biometric categories.
LLaVA and
Safe-LLaVA across different biometric categories.
As shown, LLaVA dataset includes explicit mentions of sensitive attributes like gender, age, race, and weight.
In contrast, Safe-LLaVA replaces or refuses such content to protect privacy while retaining the overall meaning of the response.3.1.1 Biometric Information Removal Pipeline
We formalize the dataset as a collection of image-text pairs , where is a question or prompt and is its corresponding textual response. The question can either explicitly inquire about biometric attributes, denoted as , or be unrelated to biometric information, denoted as . Similarly, the response can contain biometric details, represented as , or be free from biometric attributes, denoted as . This results in three relevant types of pairs: (i) : both question and answer include biometric content, (ii) : only the answer includes biometric content, and (iii) : no biometric information is present in either. To ensure privacy compliance while preserving semantic meaning, we define a transformation function that maps each pair to a cleaned version : . The transformation handles each case as follows: